OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)? y - 2 = 0 x - 2 = 0 y - x - 2 = 0

OpenStudy (anonymous):

Find the slope of the line. The slope comes out to be 0. Therefore the equation of the line is (y-2) = 0(x-5). Equation of line comes out to be y-2=0

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

You are welcome